OpenStudy (dmezzullo):

Damien works part-time from home as a data entry professional. He enters data in forms and is paid based on the entry accuracy rate. The table below shows his compensation. A 2-column table with 'Accuracy' and 'Compensation per Form' as the column headings. Tier 1 is $14, Tier 2 is $12, and Tier 3 is $0 Based on his regular performance, the equation below can be used to determine C, the total amount Damien earns a month entering data, where x represents the total number of forms he completes that month. C = 14(0.6x) + 12(0.3x) + 0(0.1x) What is the minimum number of forms he must

OpenStudy (dmezzullo):

@ajprincess

OpenStudy (anonymous):

$90...

OpenStudy (ajprincess):

i think the last part of ur question is not complete

OpenStudy (anonymous):

its the multiple choises that isent there

OpenStudy (anonymous):

like a b c or d but its 90

OpenStudy (dmezzullo):

Damien works part-time from home as a data entry professional. He enters data in forms and is paid based on the entry accuracy rate. The table below shows his compensation. A 2-column table with 'Accuracy' and 'Compensation per Form' as the column headings. Tier 1 is $14, Tier 2 is $12, and Tier 3 is $0 Based on his regular performance, the equation below can be used to determine C, the total amount Damien earns a month entering data, where x represents the total number of forms he completes that month. C = 14(0.6x) + 12(0.3x) + 0(0.1x) What is the minimum number of forms he must complete in a month to earn a total of at least $1,080? 90 91 1068 1092

OpenStudy (dmezzullo):

there we go

OpenStudy (dmezzullo):

90 is wat i have

OpenStudy (anonymous):

then why did you ask the question lol

OpenStudy (dmezzullo):

to check

OpenStudy (ajprincess):

C = 14(0.6x) + 12(0.3x) + 0(0.1x) C=8.4x+3.6x C=12x 1080=12x x=1080/12 =90

OpenStudy (anonymous):

ok lol